OpenStudy (yrelhan4):

the distance from the point -i + 2j + 6k to the straight line through the point (2,3,-4) and parallel to the vector 6 i+ 3j - 4k.

OpenStudy (anonymous):

Seems like a work for the dot product to me.

OpenStudy (yrelhan4):

so i converted all of it into cartesian form. wrote the line as (x-2)/6=(y-3)/3=(z+4)/-4 then i put all of it =r. found out random points. then found out direction ratios of the perpendicular. through condition of perpendicularity found r. and found those random points. and the found the distance. the answer i am getting is wrong.

OpenStudy (anonymous):